OpenStudy (anonymous):

what is th evalue of f(x) =9^x when x=-2?

OpenStudy (anonymous):

What does a number being raised to a negative value represent?

OpenStudy (anonymous):

i have no ideao D:

OpenStudy (anonymous):

When a number is raised to negative power, that means the following: \[x^{-n} = \frac {1}{x^n}\]

OpenStudy (haseeb96):

A is the correct answer

OpenStudy (haseeb96):

f(x)=9^x since x=-2 now juts put -2 in the place of x f(-2) = 9^-2 =1/9^2 =1/81 is the correct answer

OpenStudy (anonymous):

thanks so much

OpenStudy (haseeb96):

if u feel that i helped u then click on best response